# EGraFFBench: Evaluation of Equivariant Graph Neural Network Force Fields for Atomistic Simulations

### v1

@@ Abstract
Equivariant graph neural networks force fields (EGraFFs) have shown great promise in modeling complex interactions in atomic systems by exploiting the graphs’ inherent symmetries. Recent works have led to a surge in the development of novel architectures that incorporate equivariance-based inductive biases alongside architectural innovations like graph transformers and message passing to model atomic interactions. However, thorough evaluations of these deploying EGraFFs for the downstream task of real-world atomistic simulations, are lacking. To this end, here we perform a systematic benchmarking of 6 EGraFF algorithms (NequIP, Allegro, BOTNet, MACE, Equiformer, TorchMDNet), to understand their capabilities and limitations for realistic atomistic simulations. In addition to our thorough evaluation and analysis of eight existing datasets based on the benchmarking literature, we release two new benchmark datasets, propose four new metrics, and three new challenging tasks. The new datasets and tasks evaluate the performance of EGraFF to out-of-distribution data, in terms of different crystal structures, temperatures, and new molecules. Interestingly, evaluation of the EGraFF models based on dynamic simulations reveals that having a lower error on energy or force does not guarantee stable or reliable simulation or faithful replication of the atomic structures. Moreover, no model clearly outperforms other models on all datasets and tasks. Importantly, we show that the performance of all the models on out-of-distribution datasets is unreliable, pointing to the need to develop a foundation model for force fields that can be used in real-world simulations. In summary, this work establishes a rigorous framework for evaluating machine learning force fields in the context of atomic simulations and points to open research challenges within this domain.
@@ 3 Models Studied
All EGraFFs employed in this work rely on equivariance in the graph structure. All models use a one-hot encoding of the atomic numbers as the node input and the position vector as a node or edge input. Equivariance in these models is ensured by the use of spherical harmonics along with radial basis functions. The convolution or message-passing implementation differs from model to model.
@@ 4.3.1 Energy and Forces
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | |||||||
| E | F | E | F | E | F | E | F | E | F | E | F | |
| Acetylacetone | 1.38 | 4.59 | 0.92 | 4.4 | 2.0 | 10.0 | 2.0 | 8.0 | 4.0 | 4.0 | 1.0 | 5.0 |
| 3BPA | 3.15 | 7.86 | 4.13 | 10.0 | 5.0 | 14.0 | 4.0 | 12.0 | 6.0 | 7.0 | 3.0 | 11.0 |
| Aspirin | 9.27 | 4.77 | 14.36 | 6.31 | 12.63 | 14.26 | 13.79 | 15.34 | 13.11 | 5.90 | 5.55 | 8.98 |
| Ethanol | 4.99 | 2.38 | 6.94 | 2.63 | 203.83 | 1305.88 | 209.96 | 1309.76 | 14.60 | 3.46 | 2.30 | 5.98 |
| Naphthalene | 2.66 | 5.19 | 5.82 | 5.46 | 182.55 | 917.09 | 161.74 | 836.82 | 6.80 | 4.34 | 2.60 | 4.03 |
| Salicylic Acid | 6.29 | 5.77 | 8.59 | 5.61 | 153.06 | 961.31 | 165.29 | 843.28 | 14.46 | 9.99 | 7.59 | 7.46 |
| LiPS | 165.43 | 5.04 | 31.75 | 2.46 | 28.0 | 13.0 | 30.0 | 15.0 | 83.20 | 51.10 | 67.0 | 61.0 |
| LiPS20 | 26.80 | 3.04 | 33.17 | 3.31 | 24.59 | 5.51 | 14.05 | 4.64 | 3274.93 | 57.63 | 20.47 | 57.19 |
| GeTe | 1780.951 | 244.40 | 1009.4 | 253.45 | 3034.0 | 258.0 | 2670.0 | 247.0 | 666.34 | 363.17 | 2613.0 | 371.0 |
@@ 4.3.2 Forward simulations
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | |||||||
| JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | |
| Acetylacetone | 28.24 | 24.55 | 29.63 | 22.17 | 30.61 | 26.04 | 31.07 | 22.90 | 29.86 | 21.78 | 29.34 | 22.49 |
| 3BPA | 0.82 | 6.02 | 1.13 | 7.98 | 1.07 | 7.13 | 0.98 | 8.36 | 0.92 | 6.99 | 0.87 | 7.31 |
| Aspirin | 0.61 | 61.98 | 0.38 | 44.11 | 0.39 | 43.45 | 0.48 | 50.35 | 0.44 | 69.47 | 0.49 | 53.06 |
| Ethanol | 0.85 | 56.13 | 1.54 | 72.16 | 1.07 | 30.62 | 5.95 | 102.47 | 0.51 | 83.91 | 1.02 | 58.33 |
| Naphthalene | 0.35 | 40.31 | 0.95 | 66.92 | 2.34 | 250.22 | 0.77 | 45.02 | - | - | 0.44 | 39.27 |
| Salicylic Acid | 0.72 | 31.55 | 1.15 | 53.69 | 1.59 | 52.90 | 1.83 | 51.12 | 0.90 | 33.52 | 0.52 | 38.46 |
| LiPS | 0.0 | 3.89 | 0.0 | 3.57 | 0.0 | 3.93 | 0.0 | 3.66 | 0.0 | 2.23 | 0.0 | 1.49 |
| LiPS20 | 0.001 | 14.92 | 0.001 | 18.32 | 0.001 | 17.08 | 0.001 | 17.70 | - | - | 0.006 | 41.70 |
| GeTe | 0.0 | 2.78 | 0.0 | 1.83 | 0.0 | 1.81 | 0.0 | 1.79 | - | - | 0.0 | 2.80 |
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | |||||||
| E | F | E | F | E | F | E | F | E | F | E | F | |
| Acetylacetone | 0.960 | 0.709 | 0.820 | 0.710 | 0.923 | 0.713 | 0.813 | 0.710 | 0.810 | 0.711 | 0.840 | 0.713 |
| 3BPA | 0.810 | 0.711 | 0.708 | 0.710 | 0.680 | 0.711 | 0.760 | 0.701 | 0.803 | 0.709 | 0.814 | 0.710 |
| Aspirin | 4.39 | 0.737 | 29.56 | 0.725 | 39.69 | 0.726 | 9.04 | 0.727 | 4.360 | 0.733 | 8822.62 | 0.782 |
| Ethanol | 15.96 | 0.740 | 1787.56 | 0.773 | 3.47 | 0.941 | 2.66 | 0.922 | 5.721 | 0.750 | 6.37 | 0.736 |
| Naphthalene | 21.39 | 0.734 | 384.35 | 0.739 | 5.08 | 0.979 | 2.18 | 0.979 | - | - | 2067.02 | 0.755 |
| Salicylic Acid | 3382.50 | 0.730 | 644.91 | 0.724 | 3.33 | 0.987 | 2.69 | 0.974 | 26.131 | 0.737 | 9.38 | 0.711 |
| LiPS | 87.52 | 0.711 | 97.64 | 0.710 | 100.07 | 0.712 | 100.30 | 0.765 | 78.93 | 0.718 | 160.60 | 0.712 |
| LiPS20 | 45.10 | 0.720 | 32.79 | 0.721 | 27.99 | 0.726 | 41.47 | 0.722 | - | - | 15108.75 | 0.834 |
| (14.206) | (0.043) | (8.09) | (0.040) | (8.201) | (0.039) | (8.613) | (0.039) | - | - | (0.065) | ||
| GeTe | 495.30 | 0.800 | 294.39 | 0.756 | 351.86 | 0.764 | 352.46 | 0.765 | - | - | 346.44 | 0.779 |
@@ 5 Concluding Insights
Limitations and future work: Our research clearly points to developing a foundation model trained on large datasets. Further, improved training strategies that (i) ensure the learning of gradients of energies and forces, (ii) take into account the dynamics during simulations, and (iii) reproduce the structure faithfully need to be developed. This suggests moving away from the traditional training approach only on energy and forces and rather focusing on the system’s dynamics. Further strategies combining experimentally observed structures and simulated dynamics can be devised through experiment-simulation fusion to develop reliable force fields that are faithful to both experiments and simulations.

### v2

@@ Abstract
Equivariant graph neural networks force fields (EGraFFs) have shown great promise in modeling complex interactions in atomic systems by exploiting the graphs’ inherent symmetries. Recent works have led to a surge in the development of novel architectures that incorporate equivariance-based inductive biases alongside architectural innovations like graph transformers and message passing to model atomic interactions. However, thorough evaluations of these deploying EGraFFs for the downstream task of real-world atomistic simulations, are lacking. To this end, here we perform a systematic benchmarking of 6 EGraFF algorithms (NequIP, Allegro, BOTNet, MACE, Equiformer, TorchMDNet), to understand their capabilities and limitations for realistic atomistic simulations. In addition to our thorough evaluation and analysis of eight existing datasets based on the benchmarking literature, we release two new benchmark datasets, propose four new metrics, and three challenging tasks. The new datasets and tasks evaluate the performance of EGraFF to out-of-distribution data, in terms of different crystal structures, temperatures, and new molecules. Interestingly, evaluation of the EGraFF models based on dynamic simulations reveals that having a lower error on energy or force does not guarantee stable or reliable simulation or faithful replication of the atomic structures. Moreover, no model clearly outperforms other models on all datasets and tasks. Importantly, we show that the performance of all the models on out-of-distribution datasets is unreliable, pointing to the need to develop a foundation model for force fields that can be used in real-world simulations. In summary, this work establishes a rigorous framework for evaluating machine learning force fields in the context of atomic simulations and points to open research challenges within this domain.
@@ 3 Models Studied
All EGraFFs employed in this work rely on equivariance in the graph structure. All models use a one-hot encoding of the atomic numbers as the node input and the position vector as a node or edge input. Equivariance in these models is ensured by the use of spherical harmonics along with radial basis functions. The convolution or message-passing implementation differs from model to model. Further hyperparameters details for all models are tabulated in App. A.10
@@
PaiNN (Schütt et al., 2021) is a polarizable atom interaction neural network consisting of equivariant message passing architecture that takes into account the varying polarizability of atoms in different chemical environments, allowing for a more realistic representation of molecular behavior.
DimeNeT++ (Gasteiger et al., 2020) is a directional message passing neural network where each rotationally equivariant message is associated with a direction in coordinate space.
@@ 4.3.1 Energy and Forces
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | PaiNN | DimeNET++ | |||||||||
| E | F | E | F | E | F | E | F | E | F | E | F | E | F | E | F | |
| Acetylacetone | 1.38 | 4.59 | 0.92 | 4.4 | 2.0 | 10.0 | 2.0 | 8.0 | 4.0 | 4.0 | 1.0 | 5.0 | 4.92 | 7.73 | 102.39 | 15.28 |
| 3BPA | 3.15 | 7.86 | 4.13 | 10.0 | 5.0 | 14.0 | 4.0 | 12.0 | 6.0 | 7.0 | 3.0 | 11.0 | 36.67 | 40.41 | 796.74 | 46.72 |
| Aspirin | 6.84 | 13.89 | 5.00 | 9.17 | 7.99 | 14.06 | 8.53 | 14.01 | 6.15 | 15.29 | 5.33 | 8.97 | 41.49 | 12.41 | 133.24 | 22.07 |
| Ethanol | 2.67 | 7.49 | 2.34 | 5.01 | 2.60 | 6.80 | 2.36 | 3.19 | 2.66 | 9.73 | 2.67 | 5.93 | 7.77 | 11.81 | 149.55 | 17.19 |
| Naphthalene | 5.70 | 6.20 | 5.14 | 2.64 | 6.67 | 6.07 | 6.26 | 1.98 | 3.88 | 7.01 | 2.55 | 4.03 | 10.56 | 4.07 | 175.04 | 19.65 |
| Salicylic Acid | 5.78 | 8.42 | 5.76 | 6.30 | 5.56 | 10.21 | 5.34 | 4.24 | 5.22 | 12.39 | 6.85 | 7.19 | 24.15 | 11.12 | 169.18 | 25.48 |
| LiPS | 165.43 | 5.04 | 31.75 | 2.46 | 28.0 | 13.0 | 30.0 | 15.0 | 83.20 | 51.10 | 67.0 | 61.0 | 128.80 | 112.43 | 55.22 | 42.23 |
| LiPS20 | 26.80 | 3.04 | 33.17 | 3.31 | 24.59 | 5.51 | 14.05 | 4.64 | 3274.93 | 57.63 | 20.47 | 57.19 | - | - | - | - |
| GeTe | 1780.951 | 244.40 | 1009.4 | 253.45 | 3034.0 | 258.0 | 2670.0 | 247.0 | 666.34 | 363.17 | 2613.0 | 371.0 | 884.28 | 330.05 | 51704.65 | 222.39 |
@@ 4.3.2 Forward simulations
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | PaiNN | DimeNET++ | |||||||||
| JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | JSD | WF | |
| Acetylacetone | 28.24 | 24.55 | 29.63 | 22.17 | 30.61 | 26.04 | 31.07 | 22.90 | 29.86 | 21.78 | 29.34 | 22.49 | 26.97 | 22.33 | 66.0 | 143.36 |
| 3BPA | 0.82 | 6.02 | 1.13 | 7.98 | 1.07 | 7.13 | 0.98 | 8.36 | 0.94 | 7.44 | 0.87 | 7.31 | 1.39 | 6.65 | 6.87 | 89.97 |
| Aspirin | 0.133 | 30.66 | 0.108 | 23.29 | 0.122 | 27.36 | 0.111 | 18.92 | 0.120 | 23.58 | 0.131 | 23.99 | 0.03 | 4.0 | 0.31 | 167.48 |
| Ethanol | 0.526 | 18.34 | 0.450 | 15.89 | 0.360 | 15.57 | 0.494 | 17.93 | 0.549 | 23.48 | 0.464 | 17.70 | 0.78 | 9.42 | 3.75 | 205.75 |
| Naphthalene | 0.089 | 20.96 | 0.082 | 19.44 | 0.093 | 24.65 | 0.095 | 22.89 | 0.090 | 26.72 | 0.081 | 19.25 | 0.02 | 2.52 | 0.23 | 130.40 |
| Salicylic Acid | 0.077 | 16.95 | 0.124 | 27.58 | 0.076 | 14.65 | 0.097 | 19.35 | 0.072 | 14.17 | 0.077 | 16.12 | 0.08 | 7.63 | 0.50 | 208.95 |
| LiPS | 0.0 | 3.89 | 0.0 | 3.57 | 0.0 | 3.93 | 0.0 | 3.66 | 0.0 | 1.97 | 0.0 | 1.49 | 0.0 | 0.51 | 0.0 | 28.55 |
| LiPS20 | 0.001 | 14.92 | 0.001 | 18.32 | 0.001 | 17.08 | 0.001 | 17.70 | - | - | 0.006 | 41.70 | - | - | - | - |
| GeTe | 0.0 | 2.78 | 0.0 | 2.06 | 0.0 | 2.03 | 0.0 | 2.02 | - | - | 0.0 | 2.80 | 0.0 | 2.29 | 0.0 | 16.77 |
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | PaiNN | DimeNET++ | |||||||||
| E | F | E | F | E | F | E | F | E | F | E | F | E | F | E | F | |
| Acetylacetone | 0.960 | 0.709 | 0.820 | 0.710 | 0.923 | 0.713 | 0.813 | 0.710 | 0.810 | 0.711 | 0.836 | 0.713 | 2.129 | 0.705 | 1.043 | 0.705 |
| (0.457) | (0.042) | (0.353) | (0.041) | |||||||||||||
| 3BPA | 0.810 | 0.711 | 0.729 | 0.710 | 0.680 | 0.711 | 0.760 | 0.710 | 0.803 | 0.709 | 0.814 | 0.710 | 0.893 | 0.716 | 1.92 | 0.707 |
| (0.446) | (0.031) | (0.367) | (0.032) | |||||||||||||
| Aspirin | 1.068 | 0.626 | 1.009 | 0.625 | 1.083 | 0.627 | 1.004 | 0.628 | 1.023 | 0.637 | 1.096 | 0.626 | 2.908 | 0.662 | 1.188 | 0.680 |
| (0.598) | (0.061) | (0.265) | (0.055) | |||||||||||||
| Ethanol | 3.287 | 0.684 | 3.497 | 0.686 | 3.239 | 0.698 | 3.579 | 0.690 | 3.252 | 0.698 | 3.420 | 0.686 | 4.828 | 0.687 | 3.071 | 0.708 |
| (1.133) | (0.070) | (0.719) | (0.073) | |||||||||||||
| Naphthalene | 2.45 | 0.624 | 2.305 | 0.603 | 2.524 | 0.599 | 2.59 | 0.604 | 2.593 | 0.616 | 2.700 | 0.604 | 4.071 | 0.661 | 1.778 | 0.693 |
| (0.839) | (0.061) | (0.520) | (0.059) | |||||||||||||
| Salicylic Acid | 2.135 | 0.625 | 1.955 | 0.604 | 2.042 | 0.621 | 2.14 | 0.610 | 1.996 | 0.616 | 2.146 | 0.594 | 4.107 | 0.687 | 2.15 | 0.694 |
| (0.696) | (0.056) | (36.01) | (0.058) | |||||||||||||
| LiPS | 87.52 | 0.711 | 97.64 | 0.710 | 100.07 | 0.712 | 100.30 | 0.765 | 78.93 | 0.718 | 160.60 | 0.712 | 662.431 | 0.705 | 222.94 | 0.699 |
| (0.049) | (89.605) | (0.042) | (42.777) | (0.052) | ||||||||||||
| LiPS20 | 45.10 | 0.720 | 32.79 | 0.721 | 27.99 | 0.726 | 41.47 | 0.722 | - | - | 15108.75 | 0.834 | - | - | - | - |
| (14.206) | (0.043) | (8.09) | (0.040) | (8.201) | (0.039) | (8.613) | (0.039) | - | - | (0.065) | - | - | - | - | ||
| GeTe | 495.30 | 0.800 | 294.39 | 0.756 | 351.86 | 0.764 | 352.46 | 0.765 | - | - | 346.44 | 0.779 | 175.928 | 0.77 | 3914.07 | 0.807 |
| - | - | (80.01) | (0.052) | (181.98) | (0.081) | |||||||||||
@@ 5 Concluding Insights
Limitations and future work: Our research clearly points to developing a foundation model trained on large datasets. Further, improved training strategies that (i) ensure the learning of gradients of energies and forces, (ii) take into account the dynamics during simulations, and (iii) reproduce the structure faithfully need to be developed. This suggests moving away from the traditional training approach only on energy and forces and rather focusing on the system’s dynamics. Further strategies combining experimentally observed structures and simulated dynamics can be devised through experiment-simulation fusion to develop reliable force fields that are faithful to both experiments and simulations.Another interesting aspect is the empirical evaluation of which particular architectural feature of a model helps in giving a superior performance for a given dataset or system (defined by the type of bonding, number of atoms, crystalline vs disordered, etc.). Such a detailed analysis can be a guide to designing improved architecture while also providing thumb rules toward the use of an appropriate architecture for a given system.
@@ A.8 Hardware details
A.9 Root mean square displacement plots
A.10 Hyperparameter details
The details of hyperparameters used for training each of the models are provided in the following tables. NequIP in Table 13, Allegro in Table 14, BOTNet in Table 15, MACE in Table 16, Equiformer in Table 17, ,TorchMDNet in Table 18, PaiNN in Table 20, and DimeNET++ in Table 19
| Hyper-parameter | Value or description |
| R max | 5.0 |
| Number of Layers | 6 |
| L max | 2 |
| Number of Features | 32 |
| Nonlinearity Type | Gate |
| Nonlinearity Scalars (e) | Silu |
| Nonlinearity Scalars (o) | Tanh |
| Nonlinearity Gates (e) | Silu |
| Nonlinearity Gates (o) | Tanh |
| Number of Basis | 8 |
| BesselBasis Trainable | True |
| Polynomial Cutoff | 6 |
| Invariant Layers | 3 |
| Invariant Neurons | 64 |
| Learning Rate | 0.005 |
| Batch Size | 1 |
| EMA Decay | 0.99 |
| EMA Use Num Updates | True |
| Early Stopping Patiences (Validation Loss) | 50 |
| Early Stopping Lower Bounds (LR) | 1.0e-6 |
| Early Stopping Upper Bounds (Cumulative Wall) | 5 days |
| Loss Coeffs (Forces) | 1 |
| Loss Coeffs (Total Energy) | 1 |
| Optimizer Name | Adam |
| LR Scheduler Name | ReduceLROnPlateau |
| LR Scheduler Patience | 5 |
| LR Scheduler Factor | 0.8 |
| Parameter | Value |
| R Max | 5.0 |
| PolynomialCutoff | 6 |
| L Max | 2 |
| Num Layers | 2 |
| Env Embed Multiplicity | 64 |
| Embed Initial Edge | True |
| Two Body Latent MLP Dimensions | [128, 256, 512, 1024] |
| Two Body Latent MLP Nonlinearity | Silu |
| Latent MLP Latent Dimensions | [1024, 1024, 1024] |
| Latent MLP Nonlinearity | Silu |
| Latent Resnet | True |
| Edge Eng MLP Latent Dimensions | [128] |
| Edge Eng MLP Nonlinearity | None |
| Learning Rate | 0.005 |
| Batch Size | 1 |
| Max Epochs | 10000 |
| EMA Decay | 0.99 |
| Early Stopping Patiences(Validation loss) | 50 |
| Early Stopping Lower Bounds(LR) | |
| Early Stopping Upper Bounds(Cumulative wall) | 5 days |
| Loss Coefficients(Forces) | 1 |
| Loss Coefficients(Total energy) | 1 |
| Optimizer Name | Adam |
| LR Scheduler Name | ReduceLROnPlateau |
| LR Scheduler Patience | 5 |
| LR Scheduler Factor | 0.8 |
| Hyper-parameter | Value or description |
| 5.0 | |
| Correlation order | 1 |
| Number of Radial basis | 8 |
| Numcber of Cutoff basis | 5 |
| 3 | |
| Number of Interactions | 5 |
| MLP Irreps | 16x0e |
| Hidden Irreps | 16x0e+16x1o+16x2e |
| Gate | Silu |
| {1:-13.663181292231226, 3:-216.78673811801755, | |
| 6:-1029.2809654211628, 7:-1484.1187695035828, | |
| 8:-2042.0330099956639, 15:-1537.0898574856286, | |
| 16:-1867.8202267974733} | |
| Forces weight | 10.0 |
| SWA Forces Weight | 1.0 |
| Energy Weight | 1.0 |
| SWA Energy Weight | 1000.0 |
| Virials Weight | 1.0 |
| SWA Virials Weight | 10.0 |
| Config type Weights | {”Default”:1.0} |
| optimizer | AMSGrad Adam |
| Batch Size | 5 |
| Validation Batch Size | 5 |
| Learning rate | 0.01 |
| SWA learning rate | 0.001 |
| Weight decay | 5e-7 |
| EMA | True |
| EMA Decay | 0.99 |
| Scheduler | ReduceLROnPlateau |
| LR factor | 0.8 |
| Scheduler patience | 50 |
| LR Scheduler gamma | 0.9993 |
| SWA | True |
| Max number of epochs | 1500 |
| Clip gradiants | 10.0 |
| Hyper-parameter | Value or description |
| 5.0 | |
| Correlation order | 3 |
| Number of Radial basis | 8 |
| Numcber of Cutoff basis | 5 |
| 3 | |
| Number of Interactions | 2 |
| MLP Irreps | 16x0e |
| Hidden Irreps | 16x0e+16x1o+16x2e |
| Gate | Silu |
| {1:-13.663181292231226, 3:-216.78673811801755, | |
| 6:-1029.2809654211628, 7:-1484.1187695035828, | |
| 8:-2042.0330099956639, 15:-1537.0898574856286, | |
| 16:-1867.8202267974733} | |
| Forces weight | 10.0 |
| SWA Forces Weight | 1.0 |
| Energy Weight | 1.0 |
| SWA Energy Weight | 1000.0 |
| Virials Weight | 1.0 |
| SWA Virials Weight | 10.0 |
| Config type Weights | {”Default”:1.0} |
| optimizer | AMSGrad Adam |
| Batch Size | 5 |
| Validation Batch Size | 5 |
| Learning rate | 0.01 |
| SWA learning rate | 0.001 |
| Weight decay | 5e-7 |
| EMA Decay | 0.99 |
| Scheduler | ReduceLROnPlateau |
| LR factor | 0.8 |
| Scheduler patience | 50 |
| LR Scheduler gamma | 0.9993 |
| SWA | True |
| Max number of epochs | 1500 |
| Clip gradiants | 10.0 |
| Hyper-parameters | Value or description |
| Equiformer | |
| Hyper-parameter | Value or description |
| Activation | Silu |
| Aggregation | Add |
| Attention Activation | Silu |
| Batch Size | 8 |
| Radius Cutoff Lower | 0.0 |
| Radius Cutoff Upper | 5.0 |
| Derivative | True |
| Early Stopping Patience | 300 |
| EMA Alpha Force | 1.0 |
| EMA Alpha Energy | 0.05 |
| Embedding Dimension | 128 |
| Energy Weight | 0.2 |
| Force Weight | 0.8 |
| Inference Batch Size | 64 |
| Learning Rate | 0.001 |
| Learning Rate Factor | 0.8 |
| Minimum Learning Rate | |
| Learning Rate Patience | 30 |
| Learning Rate Warmup Steps | 1000 |
| Max Number of Neighbors | 32 |
| Max Z | 100 |
| Neighbor Embedding | True |
| Number of Epochs | 5000 |
| Number of Heads | 8 |
| Number of Layers | 6 |
| Number of Nodes | 1 |
| Number of Radial basis function | 32 |
| Number of Workers | 6 |
| Output Model | Scalar |
| Precision | 32 |
| Radial basis function Type | Expnorm |
| Reduce Operation | Add |
| Train Size | 500 |
| Weight Decay | 0.0 |
| Hyper-parameters | Value |
| Hidden Channels | 128 |
| Output Embedding Channels | 256 |
| Interaction Embedding Size | 64 |
| Basis Embedding Size | 8 |
| Number of Blocks | 4 |
| Cutoff Distance | 5.0 |
| Envelope Exponent | 5 |
| Number of Radial Functions | 6 |
| Number of Spherical Functions | 7 |
| Number of Layers Before Skip | 1 |
| Number of Layers After Skip | 2 |
| Number of Output Layers | 3 |
| Regress Forces | True |
| Batch Size | 1 |
| Evaluation Batch Size | 1 |
| Number of Workers | 4 |
| Initial Learning Rate | 0.001 |
| Optimizer | Adam |
| Scheduler | ReduceLROnPlateau |
| Patience | 5 |
| Factor | 0.8 |
| Minimum Learning Rate | 0.000001 |
| Maximum Epochs | 2000 |
| Force Coefficient | 1000 |
| Energy Coefficient | 1 |
| Exponential Moving Average Decay | 0.999 |
| Gradient Clipping Threshold | 10 |
| Early Stopping Time | 604800 |
| Early Stopping Learning Rate | 0.000001 |
| Hyper-parameters | Value or desciption |
| Number of Atom Basis | 128 |
| Number of Interactions | 3 |
| Number of Radial Basis Functions | 20 |
| Cutoff Distance | 5.0 |
| Cutoff Network | ’cosine’ |
| Radial Basis Function | BesselBasis |
| Activation Function | silu |
| Maximum Atomic Number | 100 |
A.11 Literature comparison
| NequIP | Allegro | BOTNet | MACE | Equiformer | TorchMDNet | PaiNN | DimeNET++ | |||||||
| E | F | E | F | E | F | E | F | E | F | E | F | F | F | |
| Aspirin(Ours) | 6.84 | 13.89 | 5.00 | 9.17 | 7.99 | 14.06 | 8.53 | 14.01 | 6.15 | 15.29 | 5.33 | 8.97 | 12.41 | 22.07 |
| Aspirin(Liao and Smidt [2023]) | 5.7 | 8.0 | - | - | - | - | - | - | 5.3 | 7.2 | 5.3 | 11.0 | - | - |
| Aspirin(Fu et al. [2023]) | - | 2.3 | - | - | - | - | - | - | - | - | - | - | 9.2 | 10.0 |
| Aspirin(Thölke and Fabritiis [2022]) | - | 15.09 | - | - | - | - | - | - | - | - | 5.33 | 10.97 | - | – |
| Ethanol(Ours) | 2.67 | 7.49 | 2.34 | 5.01 | 2.60 | 6.80 | 2.36 | 3.19 | 2.66 | 9.73 | 2.67 | 5.93 | 11.81 | 17.19 |
| Ethanol(Liao and Smidt [2023]) | 2.2 | 3.1 | - | - | - | - | - | - | 2.2 | 3.1 | 2.3 | 4.7 | - | - |
| Ethanol(Fu et al. [2023]) | - | 1.3 | - | - | - | - | - | - | - | - | - | - | 5.0 | 4.2 |
| Ethanol(Thölke and Fabritiis [2022]) | - | 9.02 | - | - | - | - | - | - | - | - | 2.25 | 4.73 | - | - |
| Naphthalene(Ours) | 5.70 | 6.20 | 5.14 | 2.64 | 6.67 | 6.07 | 6.26 | 1.98 | 3.88 | 7.01 | 2.55 | 4.03 | 4.07 | 19.65 |
| NaphthaleneLiao and Smidt [2023] | 4.9 | 1.7 | - | - | - | - | - | - | 3.7 | 2.1 | 3.7 | 2.6 | - | - |
| Naphthalene(Fu et al. [2023]) | - | 1.10 | - | - | - | - | - | - | - | - | - | - | 3.8 | 5.7 |
| Naphthalene(Thölke and Fabritiis [2022]) | - | 4.21 | - | - | - | - | - | - | - | - | 3.69 | 2.64 | - | - |
| Salicylic Acid(Ours) | 5.78 | 8.42 | 5.76 | 6.30 | 5.56 | 10.21 | 5.34 | 4.24 | 5.22 | 12.39 | 6.85 | 7.19 | 11.12 | 25.48 |
| Salicylic acid(Liao and Smidt [2023]) | 4.6 | 3.9 | - | - | - | - | - | - | 4.5 | 4.1 | 4.0 | 5.6 | - | - |
| Salicylic acid(Fu et al. [2023]) | - | 1.6 | - | - | - | - | - | - | - | - | - | - | 6.5 | 9.6 |
| Salicylic acid(Thölke and Fabritiis [2022]) | - | 10.32 | - | - | - | - | - | - | - | - | 4.03 | 5.59 | - | - |
| LiPS(Ours) | 165.43 | 5.04 | 31.75 | 2.46 | 28.0 | 13.0 | 30.0 | 15.0 | 83.20 | 51.10 | 67.0 | 61.0 | 112.43 | 42.23 |
| LiPS(Fu et al. [2023]) | - | 3.7 | - | - | - | - | - | - | - | - | - | - | 11.7 | 3.2 |